The car hits the wall at 43 km/h The car piles up 45 cm and the driver still goes inside the airbag 20 cm. Calculate the acceleration (deceleration) caused to the driver


1
Expert's answer
2021-09-12T19:00:21-0400

Given:

"v_i=43\:\rm km/h=11.9\: m/s"

"v_f=0\:\rm km/h=0\: m/s"

"d_1=45\:\rm cm=0.45\: m"

"d_2=20\:\rm cm=0.20\: m"


The total displacement of the driver

"d=d_1+d_2=0.45+0.20=0.65\:\rm m"

The acceleration of the driver

"a=\frac{v_f^2-v_i^2}{2d}=\frac{0^2-11.9^2}{2*0.65}=-110\:\rm m/s^2"